# Natasha Kirova    Serguei Brazovskii

### v1

@@ 4 Excitonic insulator, multi–field model
At appearance of another, symmetry breaking, order parameters, which we associate with lattice displacements $u$, the energy functional acquires additional terms $W(\Psi)\Rightarrow W(\Psi,u)=W(\Psi)+\Delta W(\Psi,u)$ where
@@
Near the BEC-EI crossover, the variable $\Psi$, hence $\rho$, is slow in comparison with $u$ which is governed by the phonon frequency $\omega_{u}$. Then it is helpful to draw the effective potentials $W^{\ast}(\rho)=W(\rho,u_{eq}(\rho))$ and $U^{\ast}(\rho)=U(\rho,u_{eq}(\rho))$ for the equilibrium $u_{eq}(\rho)$ taken after the minimization of $W(\rho,u)$ over $u$ at a given $\rho$. Fig. 5 presents the plots for $W^{\ast}(\rho)$ and $U^{\ast}(\rho)$.
Figure 5: $W^{\ast}(\rho)=W(\rho,u_{eq}(\rho))$ and $U^{\ast}(\rho)=U(\rho,u_{eq}(\rho))$ after the minimization over $u$ at a given $\rho$.
In comparison with the energy $W(\rho)$ for the generic EI in Fig.1, now the energy function $W^{\ast}(\rho)$ demonstrates a new region V at small $\rho<\rho_{c}$, where the exciton energy $E_{ex}$ is positive and it increases with $\rho$ corresponding to the repulsion of excitons. The concave energy curve indicates that there is no self-focusing at small $\rho$, contrary to the situation for generic EI (without lattice deformations), where the convex curve starts from lowest $\rho$ (see Fig.1). The landscape of $W(\rho,u)$ for the full two-field model is presented in Fig.6, to be compared with Fig.6 in ref. [37].
Figure 6: Density plot of $W(\rho,u)$. Left: $\rho_{u}<0$, Right: $\rho_{u}>0$.
At $ho_{u}<0$ only the trivial minimum $\rho=0,u=0$ does exist. At $\rho_{u}>0$ the double metastable minimum appear at $u=\pm u_{0}=\pm\sqrt{(\rho_{u}-\rho)d/f}$. At a higher $\rho>\rho_{c}$ the double minima at $(\rho,$$\pm u_{0})$ become stable while the trivial one $(\rho=0,u=0)$ becomes metastable.
4.1 Time evolution from the numerical modeling.
For the combined model with two order parameters we shall present results of modeling for only homogeneous regimes. For the subcritical pumping $\rho_{ini}<\rho_{u}<\rho_{2}$ with the unperturbed initial $u_{ini}=0$, both $\rho$ and $u$ vain at large $t$. Small oscillations of the exciton energy appear from the macroscopic quantum interference: transfers of pairs of excitons from/to the vacuum. During all the time, $\rho(t)$ decreases monotonically while $E_{ex}$ stays nearly constant close to the unperturbed value. For the supercritical pumping $\rho_{ini}>\rho_{u}$, $\rho$ rises above the initial pumping level by generation of excitons. With $A=0.01$ we observe strong unharmonic oscillations at intermediate times. Sharp lock-in transition gives rise to the new set of strong oscillations. The corresponding time dependencies are shown at Fig.7.
Figure 7: $\rho(t)$ (blue), $u(t)$ (green) $E_{ex}(t)$ (red) for a subcritical (left panel) and a supercritical (right panel pumping.
@@ 5 Effects of hot excitons
Above, we have assumed that a quasi–condensate of optically pumped excitons appears sufficiently early as a macroscopic quantum state. It is desirable to also take into account the normal, non condensed density of excitons and its (re)conversion (from)to the condensate. Indeed, temperatures in the experiments are comparable with the estimated degeneracy temperature of the BEC, and they are even higher just after the pumping pulse. The not–quite–resonance pumping also contributes to the initial incoherent density. Within this phenomenologically oriented article, we stay away from the very important question of kinetics of non-condensed particles, relying on experimental facts of fast initial equilibration typical for PIPT.
A future microscopical study can be advanced thanks to the progress in theory of equilibration in a gas of excitons and polaritons [48, 49, 16], and to the general understanding of a non equilibrium Bose gas motivated by problems in cold atoms [14, 15]. For the ideal model of the weakly interacting Bose gas, there is a fair overlap between the regime of the microscopical kinetic and the collective NLSE–based descriptions. But the price is that turbulent mixing must be taken into account in the NLSE [13, 50] and the Gross–Pitaevskii equation must be considered stochastically rather than deterministically [16]. For applications in solid state physics, the universal Bose gas mode is not much helpful as regards the BEC of excitons, because other channels of the relaxation like the emission of phonons [51] become more important than collisions of bosons.
To elucidate these effects we shall present below two types of modeling: i. stochastic equations taking into account the particles falling to the condensate as a random noise; ii. phenomenological description for a dynamical equilibrium with explicit exchange among the condensate and the incoherent condensate.
5.1 Stochastic modeling at $d=0$.
The conversion of particles from the normal to condensate reservoirs and the related decoherence can be modeled numerically for a zero-dimensional system. We need to generalize the NLSE to a stochastic Wigner- Stratonovich process. The type of the random process must be chosen to correspond to random acts of absorption of non-condensed particle into the condensate. The random function $\nu$ needs to be added to the equation for the condensate density $\rho$. That does not affect directly the equation for the phase, but the final effect of the randomness upon the phase will be the most profound as it is demonstrated by the numerical solution presented below.
@@
Fig.9 presents the plots for the phase and the density of the condensate and for the corresponding realization of random process (chosen as the exponential Brownian motion). The random increments of $\rho$ are transferred to phase kicks, hence to the decoherence. Another effect, seen at higher $T$ is the sequence of lock-in attempts with switchings among different lock-in values of the phase.
Figure 8: Plots of $\rho(t),$ $u(t)$,$E_{ex}(t)$ for a subcritical (left panel) and a supercritical (right panel) pumping.
@@ 5.2 The account for incoherent excitons and a two–fluid description
Hidden problem in experimental feasibility of the BEC-EI scenario is the transformation of the initially incoherent ensemble of pumped excitons to their coherent Bose condensate. Initially hot excitons are expected to form a conventional static ensemble of Bose particles, which occupation numbers of normal states $|k>$ are $n_{k}\sim 1$ with respect to the total large number of particles; their wave functions $\psi_{k}\varpropto\exp(ikx-E_{k}t)$. Being allowed to be quasi-static, which requires for some bottlenecks, this ensemble can be described by Boltzmann kinetics. A more complicated quasi-stationary version is a diffusive energy flow towards low energies. Under this cooling process an intermediate state of a quasi-condensate is formed where at lowest $E_{k}$ the occupation numbers start to be large $n_{k}>>1$. This part of the ensemble can be described by a quasi-classical Gross-Pitaevskii equation, still in the turbulent regime for the condensate with the kinetics for the down-flow of normal particles. At the final equilibration the occupation numbers of lowest eigenstates $n_{0}\sim N$ and they can be described by the single wave function augmented by the equilibrium Bose-Einstein distribution (with the chemical potential locked at $E_{0}$) for normal particles.
The theory is simplified by considering the normal component as a separate quasi–equilibrium reservoir, which can be characterized by the density $n$ or the chemical potential $\mu_{n}$. This approach is validated by particular spectra of polaritons, but it still needs to be justified for our system; we briefly outline its possible application below as an absolutely minimalistic description. A certain ground for the separation into two distinct, particle–exchanging reservoirs comes from suggesting a bottleneck — a minimum $E_{min}$ of the kinetic energy — where the pumped excitations are accumulated after the initial rapid cooling. It is tempting to associate $E_{min}$ with the energy of the lowest lattice mode interacting with excitons. In our case, a good candidate is the soft mode in the dip of the Kohn anomaly, which should exist as a precursor for the lattice dimerization instability, (see [52] and the references therein). That can also be the Debye frequency of the acoustical spectrum; both candidates converge to $E_{min}\sim 100\ K$.
@@
$i\hbar\partial_{t}\Psi=-\frac{\hbar^{2}}{2M}\partial_{x}^{2}\Psi+U(|\Psi|^{2}+n)\Psi+\frac{i}{2}Rn\Psi-A/2\Psi^{\ast}$
@@
The simpler space–independent Eqs. (14,15) are generalized as
@@
The results for the case of pumping to the reservoir of uncondensed excitons are presented in Fig. LABEL:fig:9,_fig.10.
Figure 9: Subcritilal pumping, $n(0)=0.27$. Left panel : Plots for $\rho(t)$, $n(t)$, $E_{ex}(t)$,$\pi\zeta(t)/100$; Right panel: Parametric polar plot $(\rho,\zeta)$
Figure 10: Supercritical pumping, $n(0)=0.28$. Left panel: Plots for $\rho(t)$, $n(t)$, $E_{ex}(t)$,$\pi\zeta/2(t)$ Right panel: Parametric polar plot $(\rho,\zeta)$.
At a subcritical pumping $n(0)=0.27$ all quantities show oscillations which are particularly pronounced in $E_{ex}(t)$. The instantaneous frequency of oscillations grows as $f(t)/\pi$. The phase decreases monotonically, correspondingly $E_{ex}(t)>0$, hence the regime of the BEC. The number of condensed particles $\rho(t)$ initially grows, being fed by the decreasing population $n(t)$ of the normal reservoir, then it passes through the maximum and vanishes at long time where $E_{ex}(t)\rightarrow E_{ex}^{0}(t)$. The polar plot shows only one trivial circulation around $\rho=0$.
For supercritical pumping $n(0)=0.28$ the phase start to decrease correspondingly $E_{ex}(t)>0$ for this initial regime of the BEC. Passing through the minimum, the phase starts to grow, correspondingly $E_{ex}(t)<0$ indicating the regime, still nonstationary, of the EI. Finally the phase is locked at the value $\zeta(t)\rightarrow\pi$. The number of condensed particles $\rho(t)$ shows an average growth superimposed by regular oscillations; it keeps growing even after the population $n(t)$ of the normal reservoir is exhausted; later the growth saturates at the equilibrium position $\rho_{st}\approx 1$ of the static EI state. $E_{ex}(t)$ shows the most irregular behavior to finally saturate at $E_{ex}(t)\rightarrow 0$ as it should be for the static EI state. The polar plot shows two circulations: first around $\rho=0$, then switching to the one around $\rho=1$.
For both cases we have presented the results very close to the dynamical pumping threshold which happens sharply between $n_{0}=0.27$ and $n_{0}=0.28$ well below the thermodynamical threshold at $q_{0}=0.36$.
@@ 6 Conclusions
The concepts and methods of pump-induced phase transitions can be applied to non-metallic systems where ubiquitous optical excitons can play a role of the pumping agents, unlike the most common practice of pumping to unbound e-h pairs. If the phase diagram of studied material shows a multitude of stable and metastable states then the pumping can achieve switching among these states or even create new static or stationary ones. In general, the excitons and the thermodynamical order parameter can be independent while mutually affecting as we have briefly outlined in Ch.2. But mostly we payed attention to a more specific class which is of a particular interest for theory and of strong experimental implications. Experimentally, this is the case of neutral-ionic transitions in donor-acceptor structures which allow for transformations by both thermodynamical means and the optical pumping. Specifically, here the excitation and the long range ordering are built from the same intermolecular electronic transfers, so that the density of pumped excitons contributes additively to the thermodynamic order. Both thermodynamical and dynamical effects can be described on the same root by viewing the ordered state as the Excitonic Insulator which appears as a macroscopic quantum state of excitons starting as their Bose condensate allowed at high pumping densities. The double nature of the ensemble of excitons leads to an intricate time evolution: the dynamical transition between number–preserved and phase–locked regimes, macroscopic quantum oscillations from interference between the Bose condensate of excitons and the ground state of the excitonic insulator. Modelling of an extended sample shows also stratification in domains of low and high densities which evolve through local dynamical phase transitions and a sequence of domains’ merging. The locally enhanced density of excitons can surpass a critical value to trigger the local persisting phase transformation, even if the mean density is below the required threshold. Beyond our phenomenological descriptions, there are challenging questions the relaxations of initial hot reservoir of excitons to the growing condensate and the establishing of the coherence. The microscopic aspects of these difficult questions have been completely resolved even in much simpler systems like cold atoms or polaritons, Here we have suggested two kinds of illustrations. One is the stochastic generalization of the macroscopic equations to emulate random absorption of incoherent excitons to the condensate. Another illustration is rate equations regulating exchange of normal and condensed reservoirs controlled by a mismatch of chemical potentials. We believe that the presented study and discussions will attract attention to still ill explored excitonic channels in the PIPT science.

### v2

@@ 4 Excitonic Insulator, Multi–Field Model
At appearance of another, symmetry breaking, order parameters which we associate with lattice displacements $u$, the energy functional acquires additional terms $W(\Psi)\Rightarrow W(\Psi,u)=W(\Psi)+\Delta W(\Psi,u)$ where
@@
Near the BEC-EI crossover, the variable $\Psi$, hence $\rho$, is slow in comparison with $u$ which is governed by the phonon frequency $\omega_{u}$. Then it is helpful to draw the effective potentials $W^{\ast}(\rho)=W(\rho,u_{eq}(\rho))$ and $U^{\ast}(\rho)=U(\rho,u_{eq}(\rho))$ for the equilibrium $u_{eq}(\rho)$ taken after the minimization of $W(\rho,u)$ over $u$ at a given $\rho$. The plots for $W^{\ast}(\rho)$ and $U^{\ast}(\rho)$ are mostly qualitatively similar to the plots for $W(\rho)$ and $U^{(}\rho)$ in Figure 1 for the generic EI. The difference is that now the energy function $W^{\ast}(\rho)$ demonstrates a new convex region at small $\rho<\rho_{u}$, where the exciton energy $E_{ex}$ is positive and it increases with $\rho$ corresponding to the repulsion of excitons. The convex energy curve indicates that now there is no self-focusing at small $\rho$, contrary to the situation for the generic EI (without lattice deformations), where the concave curve starts from lowest $\rho$ (see Figure 1). The landscape of $W(\rho,u)$ for the full two-field model is presented in Figure 5, to be compared with figure 6 in ref. [36].
Figure 5: Density plot of $W(\rho,u)$ for $\rho_{u}>0$. The double minima in $u$ are formed at the level $q\approx 0.2$.
At $\rho_{u}<0$, only the trivial minimum at $\rho=0,u=0$ does exist. At $\rho_{u}>0$ the double metastable minimum appear at $u=\pm u_{0}=\pm\sqrt{(\rho_{u}-\rho)d/f}$. At a higher $\rho_{u}>\rho_{uc}$ the double minima becomes stable while the trivial one $(\rho=0,u=0)$ persists as a metastable one.
Time Evolution from the Numerical Modeling
For the combined model with two order parameters we shall present results of modeling only for homogeneous regimes (Figure 6). For the subcritical pumping, $u(t)$ temporarily emerges from the initial $u_{ini}=0$, then at large $t$ it vanishes together with $\rho\rightarrow 0$ while $E_{ex}$ stays nearly constant close to the unperturbed value. For the supercritical pumping both $\rho$ and $u$ reach, at long $t$, the new nontrivial equilibrium with vanishing exciton energy and the locked phase. All quantities show two types of oscillations: common ones from the quantum interference and the specific ones with the frequency $\omega_{u}$ of the phonon mode in (17).
Figure 6: Plots for $\rho(t)$ – red, $u(t)$ – green, and $E_{ex}(t)$ – black. The left panels show results for a nearly subcritical pumping. The right panel shows results for a just supercritical pumping, with the added plot for the phase (in blue, rescaled).
@@ 5 Effects of Hot Excitons
Above, we have assumed that a quasi–condensate of optically pumped excitons appears sufficiently early as a macroscopic quantum state. It is desirable to take into account also the normal, non condensed density of excitons and its (re)conversion (from)to the condensate. Indeed, temperatures in the experiments are comparable with the estimated degeneracy temperature of the BEC, and they are even higher just after the pumping pulse. The not–quite–resonance pumping also contributes to the initial incoherent density. Within this phenomenologically oriented article, we stay away from the very important question of kinetics of non-condensed particles, relying on experimental facts of the fast initial equilibration typical for PIPT.
A future microscopical study can be advanced thanks to the progress in theory of equilibration in a gas of excitons and polaritons [48, 49, 16], and to the general understanding of a non equilibrium Bose gas motivated by problems in cold atoms [14, 15]. For the ideal model of the weakly interacting Bose gas, there is a fair overlap between the regime of the microscopical kinetic and the collective NLSE–based descriptions. But the price is that a turbulent mixing must be taken into account in the NLSE [13, 50] and the Gross–Pitaevskii equation must be considered stochastically rather than deterministically [16]. For applications in solid state physics, the universal Bose gas model is not quite helpful as regards the BEC of excitons, because other channels of the relaxation like the emission of phonons [51] become more important than collisions of bosons.
To elucidate these effects we shall present below two types of modeling: i. stochastic equations taking into account the particles falling to the condensate as a random noise; ii. phenomenological description for a dynamical equilibrium with explicit exchange among the condensate and the incoherent reservoir.
5.1 Stochastic Modeling at $d=0$
The conversion of particles from the normal reservoir to the condensate and the related decoherence can be modeled numerically for a zero-dimensional system. We need to generalize the NLSE to a stochastic differential equation. The type of the random process must be chosen to simulate the random acts of absorption of non-condensed particle into the condensate. The random function $\nu$ needs to be added to the equation for the condensate density $\rho$. That does not affect directly the equation for the phase, but the final effect of the randomness upon the phase will be the most profound as it is demonstrated by the numerical solution presented below.
@@ 5.1 Stochastic Modeling at $d=0$
Figure 7 presents the plots for the phase and the density of the condensate and for the corresponding realization of the random process (chosen as the exponential Brownian motion). The random increments of $\rho$ are transferred to phase kicks, hence to the decoherence. Another effect, seen at higher $T$ is the sequence of lock-in attempts with switchings among different lock-in values of the phase.
Figure 7: Plots of $\nu(t)\times 10$ – green, $\rho(t)\times 10$ – red, $\zeta(t)/\pi$ for a supercritical pumping.
@@ 5.2 The Account for Incoherent Excitons and a Two–Fluid Description
[…]
The theory is simplified by considering the normal component as a separate quasi–equilibrium reservoir, which can be characterized by the density $n$ or the chemical potential $\mu_{n}$. This approach is validated by particular spectra of polaritons, but it still needs to be justified for our system; we briefly outline its possible application below as an absolutely minimalistic description. A certain ground for the separation into two distinct, particle–exchanging reservoirs comes from suggesting a bottleneck – a minimum $E_{min}$ of the kinetic energy – where the pumped excitations are accumulated after the initial rapid cooling. It is tempting to associate $E_{min}$ with the energy of the lowest lattice mode interacting with excitons. In our case, a good candidate is the soft mode in the dip of the Kohn anomaly, which should exist as a precursor for the lattice dimerization instability, (see [52] and the references therein). That can also be the Debye frequency of the acoustical spectrum; both candidates converge to $E_{min}\sim 100$ K.
@@
$i\hbar\partial_{t}\Psi=-\frac{\hbar^{2}}{2M}\partial_{x}^{2}\Psi+U(|\Psi|^{2}+n)\Psi+\frac{i}{2}Rn\Psi-\frac{A}{2}\Psi^{\ast}$
@@
The simpler space–independent Equations (13) and (14) are generalized as
@@
The results for the case of pumping to the reservoir of uncondensed excitons have been presented above in plots of Figure 3. At the subcritical pumping $n(0)=0.27$ all quantities show oscillations which are particularly pronounced in $E_{ex}(t)$. The phase decreases monotonically, correspondingly $E_{ex}(t)>0$, hence the regime of the BEC. The concentration of condensed particles $\rho(t)$ initially grows, being fed by the decreasing population $n(t)$ of the normal reservoir, then it passes through the maximum and vanishes at long time where $E_{ex}(t)\rightarrow E_{ex}^{0}(t)$. The polar plot shows only one trivial circulation around $\rho=0$.
For the supercritical pumping $n(0)=0.28$ the phase start to decrease, correspondingly $E_{ex}(t)>0$ for this initial regime of the BEC. Passing through the minimum, the phase starts to grow, correspondingly $E_{ex}(t)<0$ indicating the regime, still nonstationary, of the EI. Finally the phase is locked at the value $\zeta(t)\rightarrow\pi$. The number of condensed particles $\rho(t)$ shows an average growth superimposed by regular oscillations; it keeps growing even after the population $n(t)$ of the normal reservoir is exhausted; later the growth saturates at the equilibrium position $\rho_{st}\approx 1$ of the static EI state. $E_{ex}(t)$ shows the most irregular behavior to finally saturate at $E_{ex}(t)\rightarrow 0$ as it should be for the static EI state. The polar plot shows two circulations: first around $\rho=0$, then switching to the one around $\rho=1$.
For both cases we have presented the results very close to the dynamical pumping threshold which happens sharply between $n_{0}=0.27$ and $n_{0}=0.28$ well below the thermodynamical threshold at $\rho_{0}=0.36$.
@@ 6 Conclusions
The concepts and methods of pump-induced phase transitions can be applied to non-metallic systems where ubiquitous optical excitons can play a role of pumping agents, unlike the most common practice of pumping to unbound e-h pairs. If the phase diagram of studied material shows a multitude of stable and metastable states, then the pumping can achieve switching among these states or even create new static or stationary ones. In general, the excitons and the thermodynamical order parameter can be independent while mutually affecting as we have briefly outlined in Ch.2. But mostly we payed attention to a more specific class which is of a particular interest for theory and of strong experimental implications. Experimentally, this is the case of neutral-ionic transitions in donor-acceptor structures which allow for transformations by both thermodynamical means and the optical pumping. Specifically, here the excitation and the long range ordering are built from the same intermolecular electronic transfers, so that the density of pumped excitons contributes additively to the thermodynamic order. Both thermodynamical and dynamical effects can be described on the same root by viewing the ordered state as the Excitonic Insulator which appears as a macroscopic quantum state of excitons starting as their Bose condensate allowed at high pumping densities. The double nature of the ensemble of excitons leads to an intricate time evolution: the dynamical transition between number–preserved and phase–locked regimes, macroscopic quantum oscillations from interference between the Bose condensate of excitons and the ground state of the excitonic insulator. Modeling of an extended sample shows also stratification in domains of low and high densities which evolve through local dynamical phase transitions and a sequence of domains’ merging. The locally enhanced density of excitons can surpass a critical value to trigger the local persisting phase transformation, even if the mean density is below the required threshold.
Beyond our phenomenological descriptions, there are challenging questions of the relaxations of initial hot reservoir of excitons to the growing condensate and of the establishing of the coherence. The microscopic aspects of these difficult questions have not been completely resolved even in much simpler systems like cold atoms or polaritons. Here we have suggested two kinds of illustrations. One is the stochastic generalization of the macroscopic equations to emulate random absorption of incoherent excitons to the condensate. Another illustration is the rate equation regulating exchange of normal and condensed reservoirs controlled by a mismatch of chemical potentials.
We believe that the presented study and discussions will attract attention to still ill explored excitonic routes in the PIPT science.